# Johannes Matzat        Aiko Schmeißer

### v4

@@ Abstract
Labor unions influence economic outcomes not only through bargaining with employers over work contracts but also via political activities that can profoundly shape political systems. In unionized workplaces, they may mobilize and change the ideological positions of both unionizing workers and their non-unionizing management. In this paper, we analyze the workplace-level impact of unionization on workers’ and managers’ political campaign contributions in the United States from 1980 to 2016. To do so, we link establishment-level union election data with transaction-level campaign contributions to federal and local candidates. Using a difference-in-differences design, validated through regression discontinuity tests and a novel instrumental variable approach, we find that unionization leads to a leftward shift of campaign contributions. Unionization increases the support for Democrats relative to Republicans not only among workers but also among managers, suggesting that it fosters political alignment between the two groups. Our findings are not driven by compositional changes of the workforce and are stronger in establishments where, after union certification, the employer and union agree on a collective bargaining agreement.
@@ 1 Introduction
Labor market institutions are typically evaluated by economists based on their ability to solve market failures and improve economic efficiency. Yet, the social welfare implications of labor market institutions also hinge critically on their potential ramifications for the political system (Acemoglu and Robinson, 2013). Labor unions have historically functioned as institutions dedicated to redistributing both economic and political power within society. While a long tradition of research has focused on how unions affect work contracts of their members via collective bargaining (Card, 1996; DiNardo and Lee, 2004; Farber et al., 2021; Frandsen, 2021; Knepper, 2020), unions seek to shape economic outcomes more broadly by engaging in various political activities. They are often viewed as one of the few vehicles that give political voice to workers, counterbalancing the dominant influence of business elites in politics (Acemoglu and Johnson, 2023; Rosenfeld, 2014; Schlozman, 2015). [^1] Unions have been credited with shaping key welfare and labor market policies, such as the 8-hour workday, minimum wages, safety standards, and sickness and family leave (e.g., Biden, 2021; King, 1965; Obama, 2010). They invest significant resources into political activities. In 2010, labor unions in the United States employed over 3,000 full-time political workers and spent 700 million USD on political outreach, a figure that rose to 1.8 billion USD in 2020 (NILRR, 2021; WSJ, 2012).
The greatest political leverage of U.S. labor unions likely stems from their connection with more than 14 million union members and their colleagues in unionized workplaces. After family and friends, the workplace is one of the most important arenas for social interaction and political discussion (Hertel-Fernandez, 2020), making it a particularly influential space for unions. By providing political information and training as well as facilitating communication networks between members, unions can mobilize workers and affect their ideological positions. Nonetheless, the overall political influence of unions in the workplace is far from clear. Even if unions successfully rally unionized workers around their political positions, it is unclear whether they can also persuade the firm’s management. Heightened tensions between workers and managers, who act on behalf of ownership interests, may provoke adverse reactions to labor demands. Any backlash in the political behavior of this powerful out-group may prevent unions from shifting political power toward worker interests.
@@
To start with, our approach enables us to analyze the political effects of unions in the workplace, where both unionizing workers and their non-unionizing management can be affected. Exploiting the occupational information in the campaign contribution data, we can differentiate the political responses of workers and managers, allowing us to explore within-firm dynamics that have not been studied before. We first ask how workplace unionization alters the political behavior of workers. As Kerrissey and Schofer (2013) argue, unions provide their members with political capital – they inform, engage, and mobilize members. Unions dedicate significant resources to political outreach and education, informing members on issues related to their working conditions and typically advocating Democratic Party positions (Ahlquist and Levi, 2013; Iversen and Soskice, 2015; Macdonald, 2021). Moreover, employee gatherings, voting for union officers, participation in hiring halls, and joint strike activities can improve communication networks between workers and create social experiences that turn them into more engaged citizens (Lindvall, 2013; McAdam et al., 2001; Terriquez, 2011). Through these mechanisms, unionization can change the political participation and ideological positions of their members.
@@
Our main results document that successful workplace unionization leads to significant increases in workers’ political donations, benefiting Democratic candidates. We find an increase in total campaign contributions from workers by 11% in response to unionization. This effect only shows up in the cycle of the union election, suggesting a short-term political mobilization of workers in response to a successful union campaign. Most importantly, when we examine the party composition of contributions, we find that unionization increases the percentage difference in donations from workers to Democrats versus Republicans by 12 percentage points in the six years following a union election. This result indicates a lasting shift in workers’ ideological positions towards the political left.
@@
Focusing solely on union members would overlook a critical out-group: the firm’s management, whose responses can either amplify or counteract the political influence of unions. On one hand, unionization may foster greater understanding of worker issues. Unions give workers a collective voice, presenting their preferences on a more equal footing with management (Freeman and Medoff, 1979, 1984). The establishment of clear bargaining rules may improve both the quantity and quality of communication between managers and workers (Verma, 2005). According to contact theory, this increase in cooperative interactions can enhance perspective-taking and, ultimately, lead to an alignment of ideological positions (e.g., Allport, 1954). [^2] Ash et al. (2019) find that granting workers greater authority through entitlements in collective bargaining agreements reduces labor conflicts, as seen in fewer and less intense strikes after negative wage shocks. On the other hand, labor unions may provoke a backlash from management. Representing the interests of firm owners, managers are typically profoundly hostile to unionization. [^3] In the run-up to union elections, employers frequently hire anti-union law firms and consultants, try to delay the election process, hold meetings in which employees are obligated to listen to the anti-unionization arguments, and – although legally restricted – threaten employees with dismissals and establishment closures (Flanagan, 2007; Freeman and Kleiner, 1990a; Kleiner, 2001; Logan, 2002; Schmitt and Zipperer, 2009). The group threat hypothesis posits that the increased bargaining power of workers may be perceived by managers as undermining their own status, power, and economic interests (e.g., Sherif et al., 1961). When unionization heightens the salience of distributional conflicts and reinforces worker-manager identities, it may increase political polarization, as individuals increasingly adopt the stereotypes associated with these identities (Bonomi et al., 2021). Thus, it is unclear ex ante whether labor unions can persuade managers to align with workers’ political positions or whether they reinforce the management’s opposition.
@@
Our results notably reveal a leftward shift in campaign contributions not only for workers but also for managers. By exploiting occupational information in the campaign contribution data, we can directly estimate the response of managers. We find that unionization increases the relative difference in managers’ donations to Democratic versus Republican candidates by 20 percentage points, without affecting their total spending. These results suggest that, instead of increasing tensions, unionization fosters a convergence in ideological positions between workers and their management.
@@
Combining establishment-level political outcomes with variation in unionization after union elections also provides plausible identification strategies to identify the causal impact of unionization on employees’ political behavior. Since we only consider establishments with union elections, i.e., where workers have shown an interest in unionization, our sample can be expected to be more similar than a random sample of establishments. Within this sample, we compare campaign contributions between establishments where workers voted for and against unionization by estimating a stacked DiD model. The stacking approach addresses weighting issues that arise under staggered treatment timing and heterogeneous treatment effects (Goodman-Bacon, 2021). [^4] Our results are robust to employing alternative DiD estimators for settings with staggered treatment timing (Borusyak et al., forthcoming; Callaway and Sant’Anna, 2021).
@@
We validate the underlying parallel trends assumption by complementing the DiD framework with tests originating from a regression discontinuity design (RDD) and with a novel instrumental variables (IV) approach. First, we test whether changes in outcomes are correlated with the pro-union vote share among establishments that lost the union election. Since the treatment status changes discontinuously at the 50% threshold, there should be no differential trends across establishments with different vote shares below 50%. Indeed, we do not find any evidence for differential changes across different vote shares, which helps us to rule out the possibility that any sizeable confounding factors correlated with the pro-union vote share and the timing of the election drive the results. Second, we restrict the sample to establishments with increasingly close elections, which are more likely to follow similar trends in contribution patterns. Our results are robust to a wide range of vote share bandwidths around the 50% cutoff, even when focusing on elections decided by only a 5-10% margin. Finally, we complement the DiD with arguably exogenous variation in union support, exploiting random shocks to the salience of workplace safety that are triggered by unexpected spikes in sector-level fatal work accidents shortly before the union election. The DiD-IV results confirm our main findings.
@@
The establishment-level effects of unionization on employees’ campaign contributions could be explained by changes in the contributions of individual employees or by changes in the composition of the employed workforce. To differentiate between individual-level and compositional effects, we use donor identifiers and track each donor’s campaign donations over time. We develop two specifications. First, we study individual-level effects by focusing on incumbent employees who donated and were employed at the establishment before the union election. We find a significant leftward shift in donations from individual workers and managers. Second, we examine compositional changes by estimating the effect on pre-election donation patterns of donors who were matched to an establishment after the election. We find no sizeable effect, indicating that the pre-existing political preferences of separated and newly hired employees do not differ across unionizing and non-unionizing establishments. In sum, these results suggest that labor unions successfully persuade both workers and their management to support Democratic candidates.
@@
Our data also allows us to move beyond party preferences by considering candidates’ ideological positions and endorsement by union organizations, as well as employees’ support of interest groups. We document considerable within-party variation in the effects on contributions to different candidates. Liberal candidates gain and conservative candidates lose, while moderate candidates are not significantly impacted on average. This suggests that our findings are not only driven by an increased signal of Democratic over Republican party affiliation but reflect shifts between candidates with clearly distinguishable ideological positions. Moreover, we differentiate candidates by whether or not they are supported by unions’ Political Action Committees (PACs). The results indicate that the partisan shift goes beyond union-endorsed candidates, reflecting a more general alignment with Democratic candidates. In addition, we show that our results extend to contributions from employees to PACs: unions mobilize workers to increase donations to union and membership PACs, while managers reduce contributions to corporate PACs. The increased support for labor and civil society interest groups from workers and the reduced support for business interest groups from managers match with the observed ideological realignment in their contributions to candidates. In contrast, we do not find any effects on the donations that the PACs of unionized firms direct to candidates. These results are inconsistent with a potential strategic explanation of changes in donations under which individuals, in particular managers, merely signal goodwill toward the interests of unions.
Finally, we show that the effects of unionization on employees’ campaign donations vary sharply with the engagement of unions, both in terms of their political activities and their ability to reach a collective bargaining contract with the employer. We first examine the role of Right-to-Work (RTW) laws, which make paying union dues for employees in unionized establishments optional. Feigenbaum et al. (2018) show that RTW laws force unions to divert scarce resources from political activities, with negative consequences for voter turnout and Democratic vote shares at the local level. Consistently, we find that the effects of unionization on support of Democratic over Republican candidates are smaller in states with RTW legislation, highlighting the importance of political activity for unions’ influence at the workplace level. Moreover, we study heterogeneity by whether a collective bargaining agreement is reached following a successful union election. Our results reveal that unionization shifts workers’ and managers’ donations to Democrats only in establishments with a bargaining contract. This suggests that unions’ political persuasion is closely tied to their ability to deliver economic benefits and foster cooperative interactions with management at the bargaining table.
@@
Our results contribute to several strands of literature. First, we complement research on the economic impacts of unions by providing insights on the political channel. Numerous studies have examined the effects of unionization on wages and compensation at U.S. establishments, typically finding limited or no wage gains with some positive effects on fringe benefits (DiNardo and Lee, 2004; Frandsen, 2021; Freeman and Kleiner, 1990b; Knepper, 2020). These modest establishment-level effects are hard to reconcile with evidence on the broader economic impact of unions. Stansbury and Summers (2020) show that declines in worker power can explain the entire decrease in the labor share of income in the U.S. over the last decades. Similarly, Western and Rosenfeld (2011) and Farber et al. (2021) find substantial negative effects of unions on income inequality, which they argue is hard to explain by income changes of union members alone, suggesting a potential link between unions and distributional legislation. [^5] Several studies point toward an important role of unions in promoting greater political representation of the working class. Sojourner (2013) shows that workers’ likelihood of serving as state legislator increases with their occupation’s unionization rate. Moreover, local union density is correlated with a more equal legislative responsiveness toward the poor versus the rich (Flavin, 2018; Becher and Stegmueller, 2021). See also Ahlquist (2017) for a review of how unions affect economic and political inequalities.
@@
Second, we speak to the literature on the political influence of unions on their members. By comparing union members to non-union members, several studies have documented a significant association with political outcomes, such as voting (Freeman, 2003; Leighley and Nagler, 2007), preferences for redistribution (Mosimann and Pontusson, 2017), and trade liberalization support (Ahlquist et al., 2014; Kim and Margalit, 2017). [^6] Union membership is also related to social attitudes more broadly, such as lower racial resentment (Frymer and Grumbach, 2021) and a stronger identification with the working class (Franko and Witko, 2023). For a recent review on labor unions in political economy, see Kaplan and Naidu (2024). We add to these studies by assessing the causal impact of unions on workers’ campaign contribution patterns. Campaign contributions are viewed as essential for candidates to win elections and have been shown to affect the selection and electoral performance of candidates (e.g., Barber, 2016b; Battaglini et al., 2024). Moreover, donors typically give to ideologically aligned candidates, such that campaign contribution patterns reveal the political ideology of donors (e.g., Bonica, 2014, 2018). Analyzing campaign contribution patterns therefore provides insights into unions’ influence on an important input into the political process and permits conclusions about shifts in political ideology.
@@
Third, we shed new light on the spread of political preferences at work by combining establishment-level union election data with an individual-level political outcome. Existing research on the political impact of unions has focused either on individual union members and their households (e.g., Freeman, 2003) or on aggregate outcomes comprising the whole county or state population (e.g., Feigenbaum et al., 2018). By focusing on the unionizing workplace, we are the first to consider within-firm dynamics, particularly the reaction of management –- an out-group that is likely indirectly affected by unionization and a key actor in political influence. Thereby, we expand on studies of politics in the workplace (Colonnelli et al., 2022; Chinoy and Koenen, 2024; Frake et al., 2024), specifically those examining spillovers in political donations between managers and workers (Babenko et al., 2020; Stuckatz, 2022) and the political effects of intergroup contact at work (Andersson and Dehdari, 2021).

### v5

@@ Abstract
Labor unions influence economic outcomes not only through bargaining with employers over work contracts but also via political activities that can profoundly shape political systems. In unionized workplaces, they may mobilize and change the ideological positions of both unionizing workers and their non-unionizing management. In this paper, we analyze the workplace-level impact of unionization on workers’ and managers’ political campaign contributions. We link establishment-level union election data with transaction-level campaign contributions to federal and local candidates in the United States. Using a difference-in-differences design, validated through regression discontinuity tests and a novel instrumental variable approach, we find that unionization leads to a leftward shift of campaign contributions. Unionization increases support for Democrats relative to Republicans not only among workers but especially among managers, suggesting that managers converge toward workers’ political preferences. The effects are stronger in settings with more cooperative union-employer interactions, such as when union elections are not contested by an unfair labor practice charge and result in a collective bargaining agreement.
@@ 1 Introduction
Labor market institutions are typically evaluated by economists based on their ability to solve market failures and improve economic efficiency. Yet, the social welfare implications of labor market institutions also hinge critically on their potential ramifications for the political system (Acemoglu2013). Labor unions have historically functioned as institutions dedicated to redistributing both economic and political power within society. While a long tradition of research has focused on how unions affect work contracts of their members via collective bargaining (Card1996; DiNardo2004; Farber2020; Frandsen2020; Knepper2020), unions seek to shape economic outcomes more broadly by engaging in various political activities. They are often viewed as one of the few vehicles that give political voice to workers, counterbalancing the dominant influence of business elites in politics (Acemoglu2023; Rosenfeld2014; Schlozman2015). [^1] Unions have been credited with shaping key welfare and labor market policies, such as the 8-hour workday, minimum wages, safety standards, and sickness and family leave (e.g., Biden2021; King1965; Obama2010). They invest significant resources into political activities. In 2010, labor unions in the United States employed over 3,000 full-time political workers and spent 700 million USD on political outreach, a figure that rose to 1.8 billion USD in 2020 (NILRR2021; WSJ2012).
The greatest political leverage of U.S. labor unions likely stems from their connection with more than 14 million union members and their colleagues in unionized workplaces. After family and friends, the workplace is one of the most important arenas for social interaction and political discussion (Hertel2020), making it a particularly influential space for unions. By providing political information and training as well as facilitating communication networks between members, unions can mobilize workers and affect their ideological positions. Nonetheless, the overall political influence of unions in the workplace is far from clear. Even if unions successfully rally unionized workers around their political positions, it is unclear whether they can also persuade the firm’s management. Heightened tensions between workers and managers, who act on behalf of ownership interests, may provoke adverse reactions to labor demands. Any backlash in the political behavior of this powerful out-group may prevent unions from shifting political power toward worker interests.
@@
To start with, our approach enables us to analyze the political effects of unions in the workplace, where both unionizing workers and their non-unionizing management can be affected. Exploiting the occupational information in the campaign contribution data, we can differentiate the political responses of workers and managers, allowing us to explore within-firm dynamics that have not been studied before. We first ask how workplace unionization alters the political behavior of workers. As Kerrissey2013 argue, unions provide their members with political capital – they inform, engage, and mobilize members. Unions dedicate significant resources to political outreach and education, informing members on issues related to their working conditions and typically advocating Democratic Party positions (Ahlquist2013; Iversen2015; Macdonald2021). [^2] Moreover, employee gatherings, voting for union officers, participation in hiring halls, and joint strike activities can improve communication networks between workers and create social experiences that turn them into more engaged citizens (Lindvall2013; McAdam2001; Terriquez2011). Through these mechanisms, unionization can change the political participation and ideological positions of their members.
@@
Our main results show that successful workplace unionization leads to notable increases in workers’ political donations, particularly in favor of Democratic candidates. We find a significant rise in total campaign contributions of workers in the cycle of the union election, suggesting a short-term mobilization effect from the union campaign. Most importantly, when we examine the party composition of contributions, we find that unionization increases the percentage difference in donations from workers to Democrats versus Republicans by 12 percentage points in the six years following a union election. [^3] The percentage interpretation of our results requires assumptions about the weighting of extensive- and intensive-margin changes in contribution amounts (Chen2024). For reasonable weights, we estimate an increase in the Democratic-Republican gap of workers’ donations between 11 and 16 percentage points. This result indicates a lasting shift in workers’ ideological positions towards the political left.
@@
Focusing solely on union members would overlook a critical out-group: the firm’s management, whose responses can either amplify or counteract the political influence of unions. On one hand, unionization may foster greater understanding of worker issues. Unions give workers a collective voice, presenting their preferences on a more equal footing with management (Freeman1970; Freeman1984). The establishment of clear bargaining rules may improve both the quantity and quality of communication between managers and workers (Verma2005). According to contact theory, this increase in cooperative interactions can enhance perspective-taking and, ultimately, lead to an alignment of ideological positions (e.g., Allport1954). On the other hand, labor unions may provoke a backlash from management. Representing the interests of firm owners, managers are typically profoundly hostile to unionization. [^4] In the run-up to union elections, employers frequently hire anti-union law firms and consultants, try to delay the election process, hold meetings in which employees are obligated to listen to the anti-unionization arguments, and – although legally restricted – threaten employees with dismissals and establishment closures (Freeman1990a; Kleiner2001; Logan2002; Schmitt2009). The group threat hypothesis posits that the increased bargaining power of workers may be perceived by managers as undermining their own status, power, and economic interests (e.g., Sherif1961). When unionization heightens the salience of distributional conflicts and reinforces worker-manager identities, it may increase political polarization, as individuals increasingly adopt the stereotypes associated with these identities (Bonomi2021). Thus, it is unclear ex ante whether labor unions can persuade managers to align with workers’ political positions or whether they reinforce the management’s opposition.
@@
Our results notably reveal a leftward shift in campaign contributions not only for workers but also for managers. By exploiting occupational information in the campaign contribution data, we can directly identify the response of managers. We find that unionization increases the relative difference in managers’ donations to Democratic versus Republican candidates by 20 percentage points, without affecting their total spending. [^5] The estimates vary between 19 and 26 percentage points for alternative extensive vs. intensive margin weights. These results suggest that, instead of increasing tensions, unionization fosters a convergence in ideological positions between workers and their management.
@@
Combining establishment-level political outcomes with variation in union elections also provides plausible identification strategies to identify the causal impact of unionization on employees’ political behavior. Since we only consider establishments with union elections, i.e., where workers have shown an interest in unionization, our sample is likely more similar than a random sample of establishments. Within this sample, we compare campaign contributions between establishments where workers voted for and against unionization in a stacked DiD model. [^6] The stacking approach addresses weighting issues that arise under staggered treatment timing and heterogeneous treatment effects (Goodman-Bacon2021). Our results are robust to employing alternative staggered DiD estimators (Borusyak2021; Callaway2021).
@@
We validate the underlying parallel trends assumption by complementing the DiD framework with tests originating from a regression discontinuity design (RDD) and with a novel instrumental variables (IV) approach. First, we test whether changes in outcomes are correlated with the pro-union vote share among establishments that lost the union election. Since the treatment status changes discontinuously at the 50% threshold, there should be no differential trends across establishments with different vote shares below 50%. Indeed, we do not find any evidence for differential changes across different vote shares, which helps us to rule out the possibility that any sizeable confounding factors correlated with the pro-union vote share and the timing of the election drive the results. Second, we restrict the sample to establishments with increasingly close elections, which are more likely to follow similar trends in contribution patterns. Our results are robust to a wide range of vote share bandwidths around the 50% cutoff, even when focusing on elections decided by only a 5-10% margin. Finally, we complement the DiD with arguably exogenous variation in union support, exploiting random shocks to the salience of workplace safety. Specifically, we use spikes in sector-level fatal work accidents shortly before the union election as an instrument for union support, and the results confirm our main findings.
@@
The establishment-level effects of unionization on employees’ campaign contributions could be explained by changes in the contributions of individual employees or by changes in the composition of the employed workforce. To differentiate between individual-level and compositional effects, we use donor identifiers and track each donor’s campaign donations over time. We study individual-level effects by focusing on incumbent employees who donated and were employed at the establishment before the union election. The results show a significant leftward shift in donations from individual workers and managers. Moreover, we examine compositional changes by estimating the effect on pre-election donation patterns of donors who were matched to an establishment after the election. We find no sizeable effect, indicating that the pre-existing political preferences of separated and newly hired employees do not differ across unionizing and non-unionizing establishments. In sum, these findings suggest that labor unions successfully persuade both workers and their management to support Democratic candidates.
@@
Our data also allows us to move beyond party preferences by considering candidates’ ideological positions and endorsement by union organizations, as well as employees’ support of interest groups. We document considerable within-party variation in the effects on contributions to different candidates. Liberal candidates gain and conservative candidates lose, while moderate candidates are not significantly impacted on average. This suggests that our findings are not only driven by an increased signal of Democratic over Republican party affiliation but reflect shifts between candidates with clearly distinguishable ideological positions. Moreover, we differentiate candidates by whether or not they are supported by unions’ Political Action Committees (PACs). The results indicate that the partisan shift goes beyond union-endorsed candidates, reflecting a more general alignment with Democratic candidates. In addition, we show that our findings extend to contributions from employees to PACs: unions mobilize workers to increase donations to union and membership PACs, while managers reduce contributions to corporate PACs. The increased support for labor and civil society interest groups from workers and the reduced support for business interest groups from managers match with the observed ideological realignment in their contributions to candidates. In contrast, we do not find any effects on the donations that the PACs of unionized firms direct to candidates. These results are inconsistent with a potential strategic explanation of changes in donations under which individuals, in particular managers, merely signal goodwill toward the interests of unions.
Finally, we show that the effects of unionization on employees’ campaign donations vary sharply with the broader labor relations environment that governs the unionization drive. We first examine the role of Right-to-Work (RTW) laws, which are typically advocated by pro-business activists to weaken unions. Feigenbaum2018 show that RTW laws diminish the political influence of unions, reducing voter turnout and Democratic vote shares at the local level. Consistent with this, we find that the establishment-level effects of unionization on support of Democratic over Republican candidates are smaller in states with RTW legislation. In addition, we reveal that in workplaces with strong managerial opposition to unionization – as indicated by an unfair labor practice charge against the employer – the political influence of unions is substantially smaller, in particular among managers. Similarly, we find that it is only when the union and the employer reach a collective bargaining agreement following the union election that unionization leads to increased support of Democratic over Republican candidates. These findings highlight the importance of cooperative labor-management interactions for unions’ ability to influence political behavior. Unions appear most effective in shaping political ideologies when they operate in environments conducive to cooperation rather than conflict.
@@
Our results contribute to several strands of literature. First, we complement research on the economic impacts of unions by providing insights on the political channel. Numerous studies have examined the effects of unionization on wages and compensation at U.S. establishments, typically finding limited or no wage gains with some positive effects on fringe benefits (DiNardo2004; Frandsen2020; Freeman1990; Knepper2020). These modest establishment-level effects are hard to reconcile with evidence on the broader economic impact of unions. Stansbury2020 show that declines in worker power can explain the entire decrease in the labor share of income in the U.S. over the last decades. Similarly, Western2011 and Farber2020 find substantial negative effects of unions on income inequality, which they argue is hard to explain by income changes of union members alone, suggesting a potential link between unions and distributional legislation. [^7] Several studies point toward an important role of unions in promoting greater political representation of the working class. Sojourner2013 shows that workers’ likelihood of serving as state legislator increases with their occupation’s unionization rate. Moreover, local union density is correlated with a more equal legislative responsiveness toward the poor versus the rich (Flavin2018; Becher2021). See also Ahlquist2017 for a review of how unions affect economic and political inequalities.
@@
Second, we speak to the literature on the political influence of unions on their members (for a recent review on labor unions in political economy, see kaplan2024between). By comparing union members to non-union members, several studies have documented a significant association with political outcomes, such as voting (Freeman2003; Leighley2007), preferences for redistribution (Mosimann2017), and trade liberalization support (Ahlquist2014; Kim2017).888yan2024 raises concerns about the causal interpretation and generalizability of these findings. Using several U.S. panel surveys, the paper finds that becoming a union member has no average effect on a range of political attitude and participation measures. In contrast to relying on self-reported changes in union membership status, we use administrative data on union elections and campaign contributions, thereby reducing measurement error issues. Moreover, our empirical design estimates the effect of first-time workplace unionization after winning a union election. This approach circumvents potential individual-level confounders correlated with becoming a union member (e.g., job transitions) and allows us to use lost union elections as a more comparable counterfactual for which we can test for parallel trends across the vote share distribution. We add to these studies by assessing the causal impact of unions on workers’ campaign contribution patterns. Campaign contributions are viewed as essential for candidates to win elections and have been shown to affect the selection and electoral performance of candidates (e.g., Barber2016a; Battaglini2024). Moreover, donors typically give to ideologically aligned candidates, such that campaign contribution patterns reveal the political ideology of donors (e.g., Bonica2014; Bonica2018). Analyzing campaign contribution patterns therefore provides insights into unions’ influence on an important input into the political process and permits conclusions about shifts in political ideology.
@@
Third, we shed new light on the spread of political preferences at work by combining establishment-level union election data with an individual-level political outcome. Existing research on the political impact of unions has focused either on individual union members and their households (e.g., Freeman2003) or on aggregate outcomes comprising the whole county or state population (e.g., Feigenbaum2018). By focusing on the unionizing workplace, we are the first to consider within-firm dynamics, particularly the reaction of management –- an out-group that is likely indirectly affected by unionization and a key actor in political influence. Thereby, we expand on studies of politics in the workplace (Colonnelli2022; Chinoy2024; Frake2024), specifically those examining spillovers in political donations between managers and workers (Babenko2020; Stuckatz2022) and the political effects of intergroup contact at work (Andersson2021). [^9] Closely related to our findings, boudreau2024 and hertelfernandez2024 also show that union leaders can play a crucial role in coordinating workers by fostering consensus around unions’ political positions and establishing norms of political participation.